9x^(5/11)-54=0

by the way x was raised to the power of 5/11
please solve. thanks for your help!

Found 2 solutions by Fombitz, Theo:
Answer by Fombitz(32388)
Answer by Theo(13342)

your equation is:
9 * x^(5/11) - 54 = 0
add 54 to both sides of this equation to get:
9 * x^(5/11) = 54
divide both sides of this equation by 9 to get:
x^(5/11) = 54 / 9 = 6
your equation is now:
x^(5/11) = 6
raise both sides of this equation by the power of (11/5).
your equation becomes:
[x^(5/11)]^11/5) = 6^(11/5)
since (x^a)^b is equal to x^(a*b), the expression on the left side of your equation becomes x
your equation is now:
x = 6^(11/5)
use your calculator to get x = 51.51488692.
that's your answer.
substitute in your original equation to confirm the answer is good.
i got 0 = 0 confirming the solution is good.